OpenStudy (anonymous):

Solve: 8x − 2(x + 1) = 7x + 8 6 −6 10 −10

OpenStudy (anonymous):

help me

OpenStudy (zpupster):

first multiply out -2(x+1) 8x - 2x-2 = 7x+8 then do the subtraction and addition of like terms 8x-2x 6x-2= 7x+8 goal is to get x by itself subtract 8 to both sides 6x-2-8=7x+8-8 6x-10 = 7x then subtract 6x from both sides )6x-6x)-10 = 7x-6x 10= 1x or just 10 = x

OpenStudy (zpupster):

sorry - 10

OpenStudy (zpupster):

(6x-6x)-10 = 7x-6x 0-10=1x -10=x

OpenStudy (anonymous):

i got it thank you very much